[Nps_daq] [EXTERNAL] Suggested EEL/108 / NPS tasks for next couple weeks

Brad Sawatzky brads at jlab.org
Sun Jul 31 12:50:58 EDT 2022


Hi all,

I'll be on vacation for the next couple weeks (Aug 2--16) so I wanted
to lay out a suggested path / bullet-list for NPS work over that time.
Please check the bullet-list below for your name and/or an 'affiliated
group'.  If I've talked to you in the last couple weeks about NPS
stuff, your name is probably on the list -- please check :-)

I'll be relying on the first name in each list to manage the
subsequent bullets -- please contact me if you have questions/
comments/concerns!

If you're a student/post-doc that is on-site *please* contact Hamlet
to support the cable bundling effort.  That is a lot of work and more
hands will be *greatly* appreciated!


Firstly, we've made a lot of progress over the last 1--2 months!
Despite all the logistical snags with our working space I am actually
quite encouraged by where we're at.  We've identified a number of
issues that need to be chased down, but that is just how this kind of
thing goes.  (And is precisely why we're doing this kind of
commissioning setup!)


The Detector and DAQ setup in EEL/108 is very close to being functional.
Basic checkout has been completed and we do see cosmics signals from the
instrumented crystals. This is great!  The readout hardware, computers,
etc are all in place and Bryan (and Ben) are working on getting the
final CODA details wrapped up.  I'm hoping we can get a pulser-driven
CODA file with sample data from 1 FADC crate ASAP.  That will give the
software folks something to chew on (replay scripts, ensure 'Hall B'
FADC firmware payloads are handled by hcana correctly, update plotting
tools with new leaf names, etc).  Then, once the noise issue is
resolved, we need to complete full checkout of the 4 instrumented
columns before the French group returns (mid-Sept).

The facilities work in EEL/108 is (largely) complete but, just to keep
things interesting, more work is going to happen in the caged area in
that space.  So, that means that we will need to clear our materials out
of that space now and find a space for them...

------------------------------------------------------
--------------------- The List (TM) ------------------
------------------------------------------------------

Hall C Techs (Walter):
- I'd like your support to move the remaining pallets of signal cables
  *and* the HV cable bundles from the cage in EEL/108 to the ESB.
- I would ask for at least 1, ideally 2 more wire baskets to help
  support that if you can locate them.  I'll check in with you Monday
  morning.

Fernando, Chris, et al:
- We'll need you guys to nail down a solution for the noise issue as
  soon as possible.   I'm expecting that we'll need to remove the
  currently installed bases and do some rework with additional grounding
  and/or EMI shielding as we discussed last week.  (It is really good
  that you can reproduce and address the noise with your setup on the
  bench -- that makes me confident that a good solution will be found!)

  The shielding/grounding changes on the 4 currently instrumented
  columns must be completed and confirmed *well* before the French group
  returns in 6 weeks to complete the remaining 950 bases.  Managing
  lead-times on purchasing/assembling components and (possibly)
  fabricating shields will be critical to maintain schedule.  Earlier
  we can get on this, the better.

  The 'overcurrent' issue associated with the instrumented LV boards
  still needs to be understood.  Take care with any in-situ LV power
  tests on the 4 boards presently installed in the NPS.

Hamlet, Vardan, Oliver, Erik, et al:
- Please continue bundling cables in the ESB
- I *strongly* encourage other NPS students/post-docs on-site to contact
  Hamlet to support this effort!

David H, Pramita, et al:
- Once an 'example' CODA file is generated from the production equipment
  (no pulses, but there will be sample-mode pedestal data), we'll need
  replay scripts and plotting scripts updated and tested.
  - David: you'll need to be in touch with Bryan for the ROC IDs, etc to
    setup the cratemap files properly (or pull them from the CODA file
    through inspection with evio2xml or similar).
  - Don't worry about unblocking

Aaron, Marc, DSG:
- Priority 1: Please focus on getting your HV (CAEN) and LV (MPOD)
  software up and running on cdaql3.  (We have workarounds, but running
  the final, production software is a priority.)
- Priority 2: User-friendly control software for the LED pulser system.
- Priority 3: Thermal readout software
    - Note: We still need to get the readout controller setup and the
        thermocouples plugged in.  This will be a task for the 2nd half
        of August.

Bryan, Ben, et al:
- Keep doing what you're doing.
- It would be great to get a running CODA environment setup an have a
  pulser-driven CODA file with sample data from 1 FADC crate (crate2)
  as soon as you can.



-- Brad

-- 
Brad Sawatzky (he/him), PhD <brads at jlab.org> -<>- Jefferson Lab/SciComp/C111
Ph: 757-269-5947  -<>-  Fax: 757-269-5235  -<>- Pager: brads-page at jlab.org
The most exciting phrase to hear in science, the one that heralds new
  discoveries, is not "Eureka!" but "That's funny..."   -- Isaac Asimov


More information about the Nps_daq mailing list